OpenStudy (surana):

What is the equation of the line which includes points (0,3) and (4,5)?

Nnesha (nnesha):

replace variables by order pair (x , y) first number in order pair is x 2nd one is y if both side equal then that order pair is ur right answer :-) remember: you should substitute both order pair values and both one should give you the equal sides

OpenStudy (surana):

I had 2 1/2. I don't think it's correct, though.

Nnesha (nnesha):

where are the equations ?? and plz show your work :-)

Nnesha (nnesha):

wait do you have the equations ??

Nnesha (nnesha):

okay got it first you have to find slope :)

OpenStudy (surana):

I put it as 0 - 4 / 3 -5.

OpenStudy (surana):

It's a bit longer, I have to find the paper first.

Nnesha (nnesha):

wait what is the formula to find slope ?

Nnesha (nnesha):

\[\huge\rm m = \frac{ y_2 - y_1 }{ x_2 - x_1}\] x should be at the denominator and y at the numerator

OpenStudy (surana):

Okay, so it should be a bit different.

Nnesha (nnesha):

just flip the fraction

OpenStudy (surana):

So 5-3 / 4 - 0?

Nnesha (nnesha):

there you go

Nnesha (nnesha):

solve that

OpenStudy (surana):

And now I have to divide it?

Nnesha (nnesha):

bec its slope which is \[\rm \frac{ rise }{ run }\] so keep that in fraction form

OpenStudy (surana):

2/4 was what I got.

Nnesha (nnesha):

yep reduce that fraction

OpenStudy (surana):

So the answer is 1/2?

Nnesha (nnesha):

yep not the answer just slope is 1/2 what's slope intercept form equation ?? do you know ?

OpenStudy (surana):

Could you refresh my memory, please?

Nnesha (nnesha):

okay \[\huge\rm y = mx +b\]slope intercept form equation where m is slope nad b is y-intercept y intercept on y -axis when x =0 when x =0 y = something <---this would be yoru y-intercept

OpenStudy (surana):

So I should put 1 into Y and 2 into x.

Nnesha (nnesha):

m is slope which is always with x m = 1/2 so just replace m by 1/2

OpenStudy (surana):

Oh, so it should be Y = 1/2x + B, then.

Nnesha (nnesha):

yep that's right now you can use one of your order pair where (x,y) first one is x and 2nd number is y plug that into the equation to find b(y-intercept ) but for this question you don't have to solve bec like i said y-intercept when x = 0

OpenStudy (surana):

Uh...so which would be the best choice, 0,3 or 4,5?

Nnesha (nnesha):

when x =0 y = ??

OpenStudy (surana):

Y is three.

Nnesha (nnesha):

so b =3

OpenStudy (surana):

Y = 1/2x + 3.

OpenStudy (surana):

Although I don't think it's finished.

Nnesha (nnesha):

yep done!

Nnesha (nnesha):

done! that is equation in slope intercept form

Nnesha (nnesha):

the*

OpenStudy (surana):

That's it? That's the answer?

OpenStudy (surana):

I didn't know it was so simple.

Nnesha (nnesha):

yep simple stuff lo_Oks hard

OpenStudy (surana):

So my answer is going to be Y= 1/2 x + 3.

OpenStudy (surana):

Thank you. I'll close this one now.

Nnesha (nnesha):

yep that's right :)